OpenStudy (anonymous):

State the various transformations applied to the base function f(x)=√x to obtain a graph of the function g(x)= √x+5 -4

OpenStudy (solomonzelman):

\(\large\color{ blue }{\large {\bbox[5pt, lightyellow ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline \texttt{Shifts} ~~~\tt from~~~ {f(x)~~~\tt to~~~g(x)}&~\tt{c~~~units~~~~} \\ \hline \\f(x)= \sqrt[ ]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[ ]{x \normalsize\color{red }{ -~\rm{c}} } &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= \sqrt[ ]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[ ]{x \normalsize\color{red}{ +~\rm{c}} } &~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= \sqrt[ ]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[ ]{x} \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= \sqrt[ ]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[ ]{x} \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

\(\large\color{black}{ g(x)=\sqrt{x+5}-4 }\) is your function. here I'll label everything for you. \(\large\color{black}{ g(x)=\sqrt{x\color{red}{+5}}\color{green}{-4} }\) the \(\large\color{black}{ \color{red}{\rm red} }\) is a shift left. the \(\large\color{black}{ \color{green}{\rm green} }\) is a shift down. (refer to the chart I posted above)

OpenStudy (solomonzelman):

have any questions?

OpenStudy (anonymous):

no i understand now Thank you

OpenStudy (solomonzelman):

You are welcome!

OpenStudy (solomonzelman):

if you want I can make a bigger chart next time, lol:)